# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v2

@@ 6.2 Sparse group Lasso
For grid search, random search and TPE, we parameterize $x_{i}$ with the log-transformed variables $\rho_{i}=\log_{10}(x_{i})$ and the search range is $[-9,2]^{J+1}$. For grid search, we set $\rho_{1}=\rho_{2}=\dots=\rho_{M}$ and use a $20\times 20$ grid. Random search samples 400 points uniformly within the range. For TPE, the search space is modeled as a uniform distribution over the same range. For VF-iDCA, MEHA, IGJO, AGILS and its variants, the initial value of $x_{0}$ is set to $[1,\dots,1]$. Both $y_{0}$ and $\theta_{0}$ for MEHA, AGILS and its two variants are initialized as $[1,\dots,1]$. For VF-iDCA, the parameters are selected as $\beta_{0}=5$, $\rho=0.01$, $c=1$, and $\delta=0.1$. VF-iDCA terminates when $\max\{\lVert{z}^{k+1}-{z}^{k}\rVert/\sqrt{1+\lVert{z}^{k}\rVert^{2}},\tilde{t}^{k+1}/m\}<0.1$, where $\tilde{t}^{k+1}$ is defined in [30]. The IGJO algorithm is run with a maximum of 50 iterations and a minimum step size of $10^{-6}$. For MEHA, the parameters are searched over $\tilde{c}_{0}\in\{1,10,20\}$ and $p\in\{0.1,0.2,0.3,0.4\}$, while $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/4,1/8\}$. The optimal parameters that minimize the validation error are selected as $\tilde{c}_{0}=20$, $p=0.1$, $\tilde{\gamma}=\gamma$, $\tilde{\alpha}=\alpha$, $\tilde{\beta}=\beta/8$, and $\tilde{\eta}=\eta/4$. MEHA terminates when $\lVert\mathbf{z}^{k+1}-\mathbf{z}^{k}\rVert/\sqrt{1+\lVert\mathbf{z}^{k}\rVert^{2}}<0.005/m$. AGILS and its variants use the same parameter settings as in the toy example, except for the penalty parameters and the following constants such as $L_{F_{x}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $L_{F_{y}}=\lambda_{\text{max}}(A_{\text{val}}^{\top}A_{\text{val}})/n_{\text{val}}$, $L_{f_{y}}=\lambda_{\text{max}}(A_{\text{tr}}^{\top}A_{\text{tr}})/n_{\text{tr}}$, $\rho_{g_{1}}=1$, $\rho_{g_{2}}=m$. The penalty parameters are set as $p_{0}=6$, $\varrho_{p}=0.01$. AGILS and its two variants terminate when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<0.005/m$, $z^{k}:=(x^{k},y^{k})$ and $t^{k+1}<0.1$. Proximal gradient method is used to compute $\theta^{k}$ that satisfies inexact criteria (17) and (20). Following the same algorithmic structure, AGILS and its variants adopt a decaying absolute tolerance $s_{k}=5/(k+1)^{1.05}$ and a relative criterion with parameter $\tau_{k}=10/(k+1)^{0.2}$.
@@
The experiments are repeated 20 times with $n_{\text{tr}}=200$, $n_{\text{val}}=200$, $n_{\text{test}}=200$, and $m=300$. Results are reported in Table 3 and Figure 3 correspond to the mean values computed over 20 independent runs, with standard deviations shown in parentheses. We report the validation error (“Val. Err.”) and test error (“Test. Err.”), computed as $\sum_{i\in I_{\text{val}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{val}}|$ and $\sum_{i\in I_{\text{test}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{test}}|$, respectively. Both errors are evaluated using the sparse group Lasso estimator, which is obtained by solving the lower-level problem with the hyperparameter value $x^{k}$ produced by the algorithms. Additionally, we report a second type of test error, referred to as the infeasible test error (“Test. Err. Infeas.”) $\sum_{i\in I_{\text{test}}}\left|b_{i}-(y^{k})^{\top}a_{i}\right|^{2}/|I_{\text{test}}|$, where the lower-level variable values are taken directly from the iterates generated by VF-iDCA, MEHA, AGILS and its variants. “Feasibility” is a scaled measure of the value function constraint violation, given by $(\varphi(x^{k},y^{k})-v(x^{k}))/|I_{\text{val}}|$ for VF-iDCA, and $(\varphi(x^{k},y^{k})-v_{\gamma}(x^{k},y^{k}))/|I_{\text{val}}|$ for MEHA, AGILS and its two variants. Note that this post-processing step is not part of the algorithm and is excluded from runtime.
@@
Figure 3: Comparison of different methods on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$. Left: Val. Err., Right: Test. Err. Infeas.
As shown in Table 3, our proposed method, AGILS, outperforms the others in terms of validation error while requiring the least computation time. The validation error, corresponding to the value of the upper-level objective at a feasible point derived from the hyperparameter value of the iterates, can be seen as an indicator of optimization performance. AGILS also achieves competitive test errors. Notably, both variants, AGILS_A and AGILS_R, exhibit similar performance to AGILS in terms of both validation and test errors. Interestingly, although VF-iDCA finds a parameter $y$ that achieves the lowest test error in the infeasible setting, the corresponding constraint violation is large. This suggests that VF-iDCA may possess good generalization performance, even when feasibility is not strictly enforced. MEHA performs similarly to AGILS, but slightly less effective. Moreover, MEHA’s performance heavily relies on proper selection of algorithmic parameters, and selecting inappropriate parameters can significantly degrade its effectiveness. Figure 3 illustrates that AGILS has fast convergence during the initial iterations. The average inner-to-outer iteration ratios for AGILS and its variants are approximately 2. This indicates that the two inexactness strategies yield similar computational efficiency.
@@
Similarly, during all repeated runs of the sparse group Lasso experiments, the feasibility correction procedure, introduced to enforce feasibility, was never triggered indicating that infeasibility was rarely encountered.
@@
$n_{\text{tr}}=1000,\ n_{\text{val}}=1000$  $\ n_{\text{test}}=1000,\ m=1500$ 24.48(3.21) 112.94(5.85) 116.72(9.49) 114.23(9.05) 0.00(0.00)
$n_{\text{tr}}=3000,\ n_{\text{val}}=3000$  $\ n_{\text{test}}=3000,\ m=4500$ 133.71(11.18), 106.89(2.91) 112.11(3.06) 109.93(2.69) 0.00(0.00)
$n_{\text{tr}}=5000,\ n_{\text{val}}=5000$  $\ n_{\text{test}}=5000,\ m=7500$ 445.44(42.99) 107.94(2.45) 106.64(1.50) 105.28(1.51) 0.00(0.00)
$n_{\text{tr}}=7000,\ n_{\text{val}}=7000$  $\ n_{\text{test}}=7000,\ m=10500$ 787.01(74.59) 105.71(1.80) 105.91(2.15) 104.72(2.14) 0.00(0.00)

### v3

@@ 6.2 Sparse group Lasso
For grid search, random search and TPE, the search space is parameterized by $\rho_{i}=\log_{10}(x_{i})$ over the range $9,2]^{J+1}$. Grid Search uses a $20\times 20$ grid with the constraint $\rho_{1}=\rho_{2}=\dots=\rho_{M}$; Random Search samples 400 points uniformly; and TPE models the search space as a uniform distribution. All iterative methods (VF-iDCA, MEHA, IGJO, AGILS and its variants) share the initial point $x_{0}=[1,\dots,1]$. For MEHA and the AGILS family, we also set $y_{0}=\theta_{0}=[1,\dots,1]$. VF-iDCA is configured with parameters $\beta_{0}=5$, $\rho=0.01$, $c=1$, and $\delta=0.1$, and it terminates when $\max\{\lVert{z}^{k+1}-{z}^{k}\rVert/\sqrt{1+\lVert{z}^{k}\rVert^{2}},\tilde{t}^{k+1}/m\}<0.1$, where $\tilde{t}^{k+1}$ is defined in [30]. The IGJO algorithm is run with a maximum of 50 iterations and a minimum step size of $10^{-6}$. For MEHA, the parameters are searched over $\tilde{c}_{0}\in\{1,10,20\}$ and $p\in\{0.1,0.2,0.3,0.4\}$, while $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters $\alpha$}, $\beta$}, $\eta$}, and $\gamma$} using scales $\{1,1/2,1/4,1/8\}$. The optimal parameters that minimize the validation error are selected as $\tilde{c}_{0}=20$, $p=0.1$, $\tilde{\gamma}=\gamma$, $\tilde{\alpha}=\alpha$, $\tilde{\beta}=\beta/8$, and $\tilde{\eta}=\eta/4$. MEHA terminates when $\lVert\mathbf{z}^{k+1}-\mathbf{z}^{k}\rVert/\sqrt{1+\lVert\mathbf{z}^{k}\rVert^{2}}<0.005/m$. AGILS and its variants use the same parameter settings as in the toy example, except for the penalty parameters and the following constants such as $L_{F_{x}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $L_{F_{y}}=\lambda_{\text{max}}(A_{\text{val}}^{\top}A_{\text{val}})/n_{\text{val}}$, $L_{f_{y}}=\lambda_{\text{max}}(A_{\text{tr}}^{\top}A_{\text{tr}})/n_{\text{tr}}$, $\rho_{g_{1}}=1$, $\rho_{g_{2}}=m$. The penalty parameters are set as $p_{0}=6$, $\varrho_{p}=0.01$. AGILS and its two variants terminate when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<0.005/m$, $z^{k}:=(x^{k},y^{k})$ and $t^{k+1}<0.1$. Proximal gradient method is used to compute $\theta^{k}$ that satisfies inexact criteria (17) and (20). The inexactness parameters: $s_{k}=5/(k+1)^{1.05}$ and $\tau_{k}=10/(k+1)^{0.2}$.
@@
The experiments are repeated 20 times with $n_{\text{tr}}=200$, $n_{\text{val}}=200$, $n_{\text{test}}=200$, and $m=300$. Results are reported in Table 3 correspond to the mean values computed over 20 independent runs, with standard deviations shown in parentheses. We report the validation error (“Val. Err.”) and test error (“Test. Err.”), computed as $\sum_{i\in I_{\text{val}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{val}}|$ and $\sum_{i\in I_{\text{test}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{test}}|$, respectively. Both errors are evaluated using the sparse group Lasso estimator, which is obtained by solving the lower-level problem with the hyperparameter value $x^{k}$ produced by the algorithms. Additionally, we report a second type of test error, referred to as the infeasible test error (“Test. Err. Infeas.”) $\sum_{i\in I_{\text{test}}}\left|b_{i}-(y^{k})^{\top}a_{i}\right|^{2}/|I_{\text{test}}|$, where the lower-level variable values are taken directly from the iterates generated by VF-iDCA, MEHA, AGILS and its variants. “Feasibility” is a scaled measure of the value function constraint violation, given by $(\varphi(x^{k},y^{k})-v(x^{k}))/|I_{\text{val}}|$ for VF-iDCA, and $(\varphi(x^{k},y^{k})-v_{\gamma}(x^{k},y^{k}))/|I_{\text{val}}|$ for MEHA, AGILS and its two variants. Note that this post-processing step is not part of the algorithm and is excluded from runtime.
@@
As shown in Table 3, our proposed method, AGILS, outperforms the others in terms of validation error while requiring the least computation time. The validation error, corresponding to the value of the upper-level objective at a feasible point derived from the hyperparameter value of the iterates, can be seen as an indicator of optimization performance. AGILS also achieves competitive test errors. Notably, both variants, AGILS_A and AGILS_R, exhibit similar performance to AGILS in terms of both validation and test errors. Interestingly, although VF-iDCA finds a parameter $y$ that achieves the lowest test error in the infeasible setting, the corresponding constraint violation is large. This suggests that VF-iDCA may possess good generalization performance, even when feasibility is not strictly enforced. MEHA performs similarly to AGILS, but slightly less effective. Moreover, MEHA’s performance heavily relies on proper selection of algorithmic parameters, and selecting inappropriate parameters can significantly degrade its effectiveness. The average inner-to-outer iteration ratios for AGILS and its variants are approximately 2.
@@
During all repeated runs of the sparse group Lasso experiments, the feasibility correction procedure, introduced to enforce feasibility, was never triggered indicating that infeasibility was rarely encountered.
@@
$n_{\text{tr}}=1000,\ n_{\text{val}}=1000$
 $n_{\text{test}}=1000,\ m=1500$
24.48(3.21) 112.94(5.85) 116.72(9.49) 114.23(9.05) 0.00(0.00)
$n_{\text{tr}}=3000,\ n_{\text{val}}=3000$
 $n_{\text{test}}=3000,\ m=4500$
133.71(11.18), 106.89(2.91) 112.11(3.06) 109.93(2.69) 0.00(0.00)
$n_{\text{tr}}=5000,\ n_{\text{val}}=5000$
 $n_{\text{test}}=5000,\ m=7500$
445.44(42.99) 107.94(2.45) 106.64(1.50) 105.28(1.51) 0.00(0.00)
$n_{\text{tr}}=7000,\ n_{\text{val}}=7000$
 $n_{\text{test}}=7000,\ m=10500$
787.01(74.59) 105.71(1.80) 105.91(2.15) 104.72(2.14) 0.00(0.00)